Solve equation: (please check to see if my answer is right- thank you so much!)
21=(4-5x)-(1-x)=4x-2(x-3)
Here is my work/answer:
21+4-5x-1-x=4x-2x+6
4-5x-1=5x-2x+27
-1=-2x+31
30=-2x
x=-15
: Solve equation: (please check to see if my answer is right- thank you so much!)
21=(4-5x)-(1-x)=4x-2(x-3)
Here is my work/answer:
21+4-5x-1-x=4x-2x+6
4-5x-1=5x-2x+27
-1=-2x+31
30=-2x
x=-15

Answer by checkley77(3848)

21=(4-5x)-(1-x)=4x-2(x-3) IF the first = sign is a + sign?????
Here is my work/answer:
21+4-5x-1-x=4x-2x+6
4-5x-1=5x-2x+27
-1=-2x+31
30=-2x
x=-15
---------------------------
Proof:
21+(4-5[-15])-(1-[-15])=4*-15-2(-15-3)
21+(4+75)-(1+15)=-60-2*-18
21+79-16=-60+36
84=-24 Obviously -15 isn't the correct answer.
-----------------------------------------------
21+4-5x+1{+x}=4x-2x+6
-5x+x-4x+2x=6-21-4-1
-6x=-20
x=-20/-6
x=10/3 ans.
Proof:
21+4-5*10/3+1+10/3=4*10/3-2*10/3+6
25-50/3+1+10/3=40/3-20/3+5
-40/3+26=20/3+6
-40/3-20/3=6-26
-60/3=-20
-20=-20
Question 169737: Solve equation: (please check to see if my answer is right- thank you so much!)
21=(4-5x)-(1-x)=4x-2(x-3)
Here is my work/answer:
21+4-5x-1-x=4x-2x+6
4-5x-1=5x-2x+27
-1=-2x+31
30=-2x
x=-15
: Solve equation: (please check to see if my answer is right- thank you so much!)
21=(4-5x)-(1-x)=4x-2(x-3)
Here is my work/answer:
21+4-5x-1-x=4x-2x+6
4-5x-1=5x-2x+27
-1=-2x+31
30=-2x
x=-15

Answer by Alan3354(1926)

21=(4-5x)-(1-x)=4x-2(x-3) Too many equal signs
Here is my work/answer:
21+4-5x-1-x=4x-2x+6 That's better
4-5x-1=5x-2x+27
-1=-2x+31
30=-2x
x=-15
----------------
21+4-5x- 1+x =4x-2x+6 Change of sign
24-5x = 2x+6
18 = 7x
x = 18/7